OpenStudy (anonymous):

The main loop on a roller coaster has a radius of 14 m. The velocity of the coaster stays constant at 15 m/s. Find the acceleration (caused by the normal force), experienced by the passengers at the following locations. Express the answers in terms of g. a) The bottom of the roller coaster b) The top of the roller coaster c) Halfway up, as the roller coaster comes down

OpenStudy (anonymous):

Use these equations on g-forces: Since g-force=apparent weight/true weight then when a going in a circular bend on a rollar coaster: Top of loop: \[\huge g-force=\frac{\frac{v^2}{r}-g}{g}\] Bottom of Loop: \[\huge g-force=\frac{\frac{v^2}{r}+g}{g}\]

OpenStudy (anonymous):

Sub in your values into the equation to find your answer. and g in this question is always 9.8 which is the acceleration due to Earth's gravity.